OpenStudy (anonymous):

Evaluate 1/3 + 1/3+ 1/3 with infinitesimals

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

.9999999999999999999999999999999999999999999999999999999999999999999...

OpenStudy (agreene):

\[\lim_{x \rightarrow 3} (x^{-1}) + lim_{x \rightarrow 3} (x^{-1}) +lim_{x \rightarrow 3} (x^{-1}) = 0.9999999... = 1\]

OpenStudy (anonymous):

i'm wondering how'd you write it in infinitesimal form... i think it'd be like infty/infty +1/infty

OpenStudy (agreene):

Well, an infinitesimal is something infinitely small (or sufficiently small). Back in the 1600s they would just pick a super small number to approximate it. 0.33333333333333333333333333333333333333333333333333333333333333 = 1/3 for example

OpenStudy (anonymous):

yeah or in otherwords \[\frac{1}{\infty}\]

OpenStudy (anonymous):

i'm wondering how'd you write .99999 in infinitesimal...like this? \[\frac{\infty}{\infty+\frac{1}{\infty}}\]

OpenStudy (agreene):

except... \[\frac{1}{\infty} = 0\] you can write .9999 as an infinitesimal, either just like that or some combination of recursive limits if you really want to.

OpenStudy (anonymous):

but 1/infty in infinitesimals i thought meant that 1/infty is not zero just extremely small to the point that they just ruled out 1/infty and said it was 0